please help me solve this equation
m=4/5,(2,-9)

Answer by stanbon(75887)

please help me solve this equation
m=4/5,(2,-9)
-------------------
Form: y = mx + b
m = (4/5) and y = -9 when x = 2.
Solve for "b":
-----
-9 = (4/5)2 + b
-9 = (8/5) + b
-(45/5) = (8/5) + b
b = - 53/5
----------------------
Equation:
y = (4/5)x - (53/5)
=========================
Cheers,
Stan H.